OpenStudy (anonymous):

What transformations change the graph of f(x) to the graph of g(x)? f(x)=2x g(x)=6x-3 I think you have to plug something in but I'm a bit confused...

OpenStudy (john_es):

\[h(x)=3x-3\] Then you do the composition, \[h(f(x))=3(2x)-3=6x-3=g(x)\]

OpenStudy (anonymous):

Where did you get h? I'm confused.

OpenStudy (john_es):

Well, by inspection, you can see, \[g(x)=6x-3=3(2x)-3\]And from this point you can deduce it must be composition.